OpenStudy (tfht):

a circle is inscribed in an equilateral triangle with a side of 6cm. find the area that is part of the triangle but not he circle

OpenStudy (anonymous):

 \[area~of~\triangle=\frac{ 1 }{2 }*6*3\sqrt{3}=9\sqrt{3}\] \[area ~of~circle=\pi r^2=\pi \left( \sqrt{3} \right)^2=3\pi \] required area=? (find the difference)

OpenStudy (tfht):

thanks!

OpenStudy (anonymous):

yw